# AutoAL: Automated Active Learning with Differentiable Query Strategy Search

### v1

@@ 4.2 Evaluation of Active Learning Performance
    Our method demonstrates robust performance not only on easier datasets but also on more challenging ones, such as Cifar-100 and OrganCMNIST. Cifar-100 has significantly more classes than other datasets, while OrganCMNIST has smaller data pool, with only $12,975$ images. These challenges make active learning more difficult, yet our method’s superior performance across these datasets highlights its robustness.
@@
    The labeled data vs. accuracy curves of our method are relatively smooth compared to those of other methods. The baseline methods sometimes show significant accuracy drops during certain active learning rounds. This mainly dues to harmful data selection Koh & Liang (2017) or overfitting problem. However, by selecting the optimal strategy thus selecting the informative data, AutoAL can alleviate the problem and make the curve much more smooth than baseline strategies.
@@
    Different AL strategies produce varying results across datasets. For instance, Margin Sampling underperforms on the SVHN dataset, while KMeans Sampling yields the worst performance on the OrganCMNIST and Cifar-10 datasets. This variability highlights the significance of our approach, which aims to identify and select the optimal strategy for each dataset.
4.3 Ablation Studies
@@
Our SearchNet consists of two main components: a ResNet-18 backbone and a loss prediction module for sample loss prediction  Yoo & Kweon (2019). In this ablation study, we disable one component at a time and record the results. Fig. 3 illustrates the performance, revealing the effectiveness of both ResNet-18 backbone and the loss prediction module. However, when only the loss prediction module is used to optimize the network, performance drops significantly. This aligns with our intuition, as although the loss prediction module can assist with network optimization, the main focus shifts from selecting the best AL strategy to merely minimizing the sample loss. This misalignment will result in incorrect AL strategy weight estimation (See Fig. 1).
@@
Since AutoAL is built on multiple AL strategies, we study how the size of the AL candidate pool impacts performance. We start from the Maximum Entropy as the sole candidate for the initial baseline, then expand the candidate pool to three strategies by adding Margin Sampling and Least Confidence. For the pool of five candidates, we further include KMeans and BALD. As shown in Fig. 4, for all three datasets, the single candidate variation performs the worst but still outperforms the Entropy Sampling baseline in Fig. 2. This is because AutoAL does not directly query data samples with maximum entropy but queries the data with the highest score from SearchNet, which has been optimized with the help of the loss prediction module and FitNet. For Cifar-100 and OrganCMNIST, AutoAL can perform well when there are at least three candidates in the selection pool. However, for SVHN, more AL candidates result in better performance. This indicates that the upper bound of candidate pool size varies across different datasets. In certain datasets, incorporating more active learning strategies yields better results, but this is not the case for other datasets. Additionally, pools with more candidates exhibit lower standard deviations. This demonstrates strong capability of AutoAL in selecting the best active learning strategy. It can consistently choose the optimal one and exclude the less effective options.
@@ 4.4 AL Selection Strategy
On Cifar-100 dataset, both Least Confidence and MeanSTD, two uncertainty-based measures, demonstrate improved performance in the final rounds. AutoAL assigns them higher priority to enhance overall performance. As the AL process progresses and more diverse data become available, the model attempts to develop a better understanding of the underlying structure of the target data. Uncertainty-based measures are more effective now because the major challenge switches to refining the decision boundaries. The transition from diversity to uncertainty-based strategies is consistent with the model’s evolving needs. In early rounds, the model requires diversity to build a broad understanding of the data, but once that foundation is established, its focus shifts toward uncertainty, which helps fine-tune model decision-making ability. This adaptive approach ensures that the model targets the most diverse subsets at each stage of the learning process, from broad exploration in the early rounds to fine-grained exploitation in the later rounds. By automatically adjusting the priorities of these strategies, AutoAL improves the overall accuracy and shows stronger robustness.

### v2

@@ 4.2 Evaluation of Active Learning Performance
    Our method demonstrates robust performance not only on easier datasets but also on more challenging ones, such as Cifar-100, OrganCMNIST and TinyImageNet. Cifar-100 and TinyImageNet has significantly more classes than other datasets, while OrganCMNIST has smaller data pool, with only $12,975$ images. These challenges make active learning more difficult, yet our method’s superior performance across these datasets highlights its robustness and generalizability.
@@
    The labeled data vs. accuracy curves of our method are relatively smooth compared to those of other methods. The baseline methods sometimes show significant accuracy drops during certain active learning rounds. This mainly dues to harmful data selection (Koh & Liang, 2017) or overfitting problem. However, by selecting the optimal strategy thus selecting the informative data, AutoAL can alleviate the problem and make the curve much more smooth than baseline strategies.
@@
    Different AL strategies produce varying results across datasets. For instance, Margin Sampling underperforms on the SVHN dataset, while KMeans Sampling and VAAL yield the worst performance on the OrganCMNIST and Cifar-10 datasets. This variability highlights the significance of our approach, which aims to identify and select the optimal strategy for each dataset.
@@ 4.3 Ablation Studies
Our SearchNet consists of two main components: a ResNet-18 backbone and a loss prediction module for sample loss prediction  (Yoo & Kweon, 2019). In this ablation study, we disable one component at a time and record the results. Fig. 3 illustrates the performance, revealing the effectiveness of both ResNet-18 backbone and the loss prediction module. However, when only the loss prediction module is used to optimize the network, performance drops significantly. This aligns with our intuition, as although the loss prediction module can assist with network optimization, the main focus shifts from selecting the best AL strategy to merely minimizing the sample loss. This misalignment will result in incorrect AL strategy weight estimation (See Fig. 1).
@@
Since AutoAL is built on multiple AL strategies, we study how the size of the AL candidate pool impacts performance. We start from the Maximum Entropy as the sole candidate for the initial baseline, then expand the candidate pool to three strategies by adding Margin Sampling and Least Confidence. For the pool of five candidates, we further include KMeans and BALD. As shown in Fig. 4, for all three datasets, the single candidate variation performs the worst but still outperforms the Entropy Sampling baseline in Fig. 2. This is because AutoAL does not directly query data samples with maximum entropy but queries the data with the highest score from SearchNet, which has been optimized with the help of the loss prediction module and FitNet. For Cifar-100 and OrganCMNIST, AutoAL can perform well when there are at least three candidates in the selection pool. However, for SVHN, more AL candidates result in better performance. This indicates that the upper bound of candidate pool size varies across different datasets. In certain datasets, incorporating more active learning strategies yields better results, but this is not the case for other datasets. Additionally, pools with more candidates exhibit lower standard deviations. This demonstrates AutoAL’s strong capability to select the best AL strategy. It can consistently choose the optimal one and exclude the less effective options.
@@ 4.4 AL Selection Strategy
On Cifar-100 dataset, both Least Confidence and MeanSTD, two uncertainty-based measures, demonstrate improved performance in the final rounds. AutoAL assigns them higher priority to enhance overall performance. As the AL process progresses and more diverse data become available, the model attempts to develop a better understanding of the underlying structure of the target data. Uncertainty-based measures are more effective now because the major challenge switches to refining the decision boundaries. The transition from diversity to uncertainty-based strategies is consistent with the model’s evolving needs. In early rounds, the model requires diversity to build a broad understanding of the data, but once that foundation is established, its focus shifts toward uncertainty, which helps fine-tune model decision-making ability. This adaptive approach ensures the model targets the most diverse subsets at each stage of learning process, from broad exploration in the early rounds to fine-grained exploitation in the later rounds. By automatically adjusting strategy priorities, AutoAL improves the overall accuracy and shows stronger robustness.
@@ 5 Conclusion
Acknowledgements
The authors acknowledge research support from Clemson University with a generous allotment of computation time on the Palmetto cluster.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning by introducing a novel automated search framework for deep active learning. Our work has many potential societal consequences, including reducing the annotation cost of training deep learning models and improving the efficiency of data annotation, etc. While our proposed method has the potential to promote decision-making in various downstream tasks (e.g., natural and medical image data selection), we acknowledge that our AutoAL has been carefully designed to mitigate potential biases and ensure fair selection.
Appendix A Appendix
Table 1: A summarization of datasets used in the experiments. The Imbalance Ratio (IR) is the ratio of the number of images in the majority class to the number of images in the minority class. $\text{IR}=1$ represents balance dataset.
DataSet Training Size Test Size Class Numbers Imbalance Ratio
Nature Images Cifar-10 50,000 10,000 10 1.0
Cifar-100 50,000 10,000 100 1.0
SVHN 73,257 26,032 10 3.0
TinyImageNet 100,000 10,000 200 1.0
Medical Images OrganCMNIST 12,975 8,216 11 5.0
PathMNIST 89,996 7,180 9 1.6
TissueMNIST 165,466 47,280 8 9.1
A.1 Complexity Analysis
Table 2: The average runtime of the entire active learning (AL) process (including training and querying), where the runtime of EntropySampling is the unit time. We separate total time cost of AutoAL into three main parts: candidate ALs querying AL score for images (AutoAL-ComputeALScore), SearchNet and FitNet updating (AutoAL-Search), and querying samples and training classification model after AutoAL search (AutoAL-TrainResNet).
CIFAR-10 CIFAR-100 SVHN TinyImageNet OrganCMNIST PathMNIST TissueMNIST
EntropySampling 1.00 1.00 1.00 1.00 1.00 1.00 1.00
MarginSampling 0.99 1.02 5.97 0.99 2.01 1.01 1.84
LeastConfidence 0.85 1.01 1.03 0.96 1.85 0.97 1.81
KMeansSampling 1.20 1.24 1.59 0.62 1.33 0.93 2.37
MeanSTD 1.46 1.18 0.89 0.23 2.07 0.85 2.35
VarRatio 1.40 1.17 0.87 0.87 1.87 0.78 2.30
BALD 1.48 1.18 0.90 0.92 2.04 1.02 1.86
BadgeSampling 1.83 4.09 2.34 8.32 1.21 1.49 2.52
LPL 2.03 0.87 0.98 0.37 1.87 1.35 1.77
ALBL 1.08 0.57 1.01 0.82 2.10 0.88 1.42
DDU 1.48 1.87 2.13 1.62 1.32 2.32 1.68
Coreset 1.68 1.42 1.07 1.34 1.25 1.47 2.05
VAAL 1.24 1.27 7.02 2.05 5.28 4.36 7.68
ENSvarR 4.21 4.52 6.58 12.27 8.33 7.63 10.46
AutoAL-ComputeALScore 2.34 2.03 2.80 3.40 3.79 3.24 1.74
AutoAL-Search 0.12 0.11 0.15 0.18 0.20 0.17 0.09
AutoAL-TrainResNet 1.57 1.36 1.88 2.28 2.55 2.18 1.15
AutoAL-Total 4.03 3.50 4.83 5.86 6.54 5.59 2.95
In this subsection, we perform additional analysis experiments to showcase AutoAL’s efficiency and generalizability. We mainly use the average running time to verify the results, and all the experiments was done on one Nvidia A100 GPU. We use the running time of EntropySampling as a benchmark and calculate other ALs running costs relative to it. We find:
  1. 1.
    Our AutoAL demonstrates its ability to generalize to large datasets, and the time cost is compatible to some of the baselines such as ENSvarR.
  2. 2.
    We separate the total time cost of our method into three main parts, and the AL strategy sampling will cost the most time and the update of SearchNet and FitNet will nearly cost none. This means AutoAL will time cost rely heavily on the candidate pool. When the candidate ALs are fast enough, AutoAL will be in low cost and complexity.
  3. 3.
    We conducted the experiment on changing the candidate pool size, and we found that when there are 3 candidates, the time cost is just 1.3x compared to EntropySampling, and our method with three candidates can also gain a high improvement on labeling accuracy. This will give real-world users flexibility in choosing the number of candidate ALs considering both accuracy and complexity.